ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 20 Krasilnikov, Alexei, Moscow Pedagogical State University, Russia Identities in metanilpotent groups The following theorem was proved jointly with R. M. Bryant. Theorem Let m, n be coprime positive integers. Then every (nilpotent of class 2 and exponent m)-by-(nilpotent of class 2 and exponent n) group has a finite basis for its identities. This solves a problem posed by L. G. Kovács and M. F. Newman in [1].

References: L. G. Kovács and M. F. Newman, Hanna Neumann’s problems on varieties of groups, Proc. II Internat. conf. theory of groups, Canberra, 1973, 417–431.